Find area enclosed between curve y=x²+2 and y=x³ , line x=1,x=2.

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
Area = integration from 1 to 2 (x3 – x2 – 2)
 = (x^4/4 – x^3 /3 – 2x) from 1 to 2
 
 = 4 - 8/3 – 4 – ¼ + 1/3 + 2
 
 = 19/3 – 83/12
 
 = 7/12